The Charvon oil company is planning to make a large investment in coal-to-liquids (CTL) gasoline. The end product will be a perfect substitute for gasoline made from petroleum, but the feedstock will be coal instead of oil. Two technologies are available to the Charvon company. The first is called indirect CTL, where the coal is gasified prior to being liquefied. The second is called direct CTL, where the coal is dissolved in a solvent, and the resulting liquid is processed into gasoline. The Charvon company has hired you as a consultant to help them decide which technology they should choose. Charvon expects to produce 1.5 million gallons of CTL gasoline in each of the next twenty years, and they can sell the gasoline for S2.5 per gallon. The capital cost of indirect CTL is S10 million and operating costs for indirect CTL (labor, fuel, and maintenance) are $400,000 per year. The capital cost of direct CTL is S12 million and operating costs for direct CTL are $300,000 per year. One problem with indirect CTL is that the coal gasification process releases large amounts of CO2 into the atmosphere. Assume that for every gallon of gasoline produced with indirect CTL, 0.02 tons of CO2 are released. You have learned that starting in Year 1 the government will implement a tax of S15 per ton of CO2, which is applied as increase in the operating cost. Assume that ifyou choose direct CTL you will not be subject to the tax on CO2. Assume capital cost for both technologies can be depreciated using straight line with project life time of 20 years, starting from year 1 to year 20. Consider income tax rate of 40%for the taxable income nd discount rate of 12% for all costs and revenues
Answer #1
CALCULATION OF NPV OF DIRECT CTL
A Initial Investment $12,000,000
B Annual Sales quantity in gallons 1500000
C Sales price per gallon $2.50
D=B*C Annual sales in dollars $3,750,000
E Annual Operating cost $300,000
F=D-E Annual Pre tax cash flow $3,450,000
G Income tax rate                                 0.40
H=F*(1-G) After tax cash flow(without considering depreciation) $                  2,070,000
I=A/20 Depreciation per year $600,000
J=I*G Depreciation tax shield $                      240,000
K=H+J Annual after tax cash flow $                  2,310,000
L Discount rate 12%
M Present Value(PV) of annual cash flow $17,254,415 (Using PV function of excel with Rate=12%,Nper=20, Pmt=2310000)
N=M-A Net Present value(NPV) $5,254,415
CALCULATION OF NPV OF INDIRECT CTL
A Cost of capital $10,000,000
B Annual sales in dollars $3,750,000
C Annual Operating cost $400,000
D=B-C Annual Pre tax cash flow $3,350,000
E Income tax rate                                 0.40
F=D8(1-E) After tax cash flow(without considering depreciation) $                  2,010,000
G=A/20 Depreciation per year $500,000
H=G*E Depreciation tax shield $                      200,000
I=F+H Annual after tax cash flow $                  2,210,000
Discount rate 12%
Present Value(PV) of annual cash flow $16,507,470 (Using PV function of excel with Rate=12%,Nper=20, Pmt=2210000)
EFFECT OF TAX ON CO2
J Amount of CO2 generated(Ton) 30000 (1500000*0.02)
K=15*J Carbon tax $                      450,000
SENSITIVITY ANALYSIS
P T=450000*P K=3350000-T L=K*(1-0.4) M=L+200000 N=PV(12%,20,-M) N-10000000
Probability of Carbon tax Rxpected NPV of Indirect CTL Expected NPV of Direct CTL Probability of $15 per ton carbon tax Expected carbon tax Pretax cash flow After tax Cash flow Annual Cash flow with depreciation Present Value of Cash flow Net Present Value
1 $4,490,721 $5,254,415 1 $                      450,000 $2,900,000 $   1,740,000 $1,940,000 $14,490,721 $4,490,721
0.9 $4,692,396 $5,254,415 0.9 $                      405,000 $2,945,000 $   1,767,000 $1,967,000 $14,692,396 $4,692,396
0.8 $4,894,071 $5,254,415 0.8 $                      360,000 $2,990,000 $   1,794,000 $1,994,000 $14,894,071 $4,894,071
0.7 $5,095,746 $5,254,415 0.7 $                      315,000 $3,035,000 $   1,821,000 $2,021,000 $15,095,746 $5,095,746
0.6215 $5,254,060 $5,254,415 0.6215 $                      279,675 $3,070,325 $   1,842,195 $2,042,195 $15,254,060 $5,254,060
0.6 $5,297,421 $5,254,415 0.6 $                      270,000 $3,080,000 $   1,848,000 $2,048,000 $15,297,421 $5,297,421
0.5 $5,499,096 $5,254,415 0.5 $                      225,000 $3,125,000 $   1,875,000 $2,075,000 $15,499,096 $5,499,096
0.4 $5,700,770 $5,254,415 0.4 $                      180,000 $3,170,000 $   1,902,000 $2,102,000 $15,700,770 $5,700,770
0.3 $5,902,445 $5,254,415 0.3 $                      135,000 $3,215,000 $   1,929,000 $2,129,000 $15,902,445 $5,902,445
0.2 $6,104,120 $5,254,415 0.2 $                        90,000 $3,260,000 $   1,956,000 $2,156,000 $16,104,120 $6,104,120
0.1 $6,305,795 $5,254,415 0.1 $                        45,000 $3,305,000 $   1,983,000 $2,183,000 $16,305,795 $6,305,795
0 $6,507,470 $5,254,415 0 $                                  -   $3,350,000 $   2,010,000 $2,210,000 $16,507,470 $6,507,470
THRESHOLD PROBABILITY OF TAX 0.6215
Probability of Carbon tax Expected NPV of Indirect CTL Expected NPV of Direct CTL
0 $6,507,470 $5,254,415
0.1 $6,305,795 $5,254,415
0.2 $6,104,120 $5,254,415
0.3 $5,902,445 $5,254,415
0.4 $5,700,770 $5,254,415
0.5 $5,499,096 $5,254,415
0.6 $5,297,421 $5,254,415
0.6215 $5,254,060 $5,254,415
0.7 $5,095,746 $5,254,415
0.8 $4,894,071 $5,254,415
0.9 $4,692,396 $5,254,415
1 $4,490,721 $5,254,415

Book1 Microsoft Excel Home Insert Page Layout Formulas Dats Review View Developer Cut Σ AutoSum) A Inherit Copy Fill 2T 스미 MBook1 Microsoft Excel Home Insert Page Layout Formulas Dats Review View Developer Cut Σ AutoSum) A Ea Copy Fill 2T 스미 MergeBook1 Microsoft Excel Dats Review V Home Insert % Cut Ea Copy Page Layout Formulas ew Developer Σ AutoSum) A Inhenit 11A A PBookl Microsoft Excel Chart Tools Home InsertPage Layout Formulas Data Review iewDeveloper Designayout Format Change Save As